OpenStudy (anonymous):

sec (–270°) A. undefined B. 1 C. 0 D. –1

OpenStudy (nurali):

sec(x) = 1/cos(x) sec(270) = 1/cos(-270) = 1/0

OpenStudy (nurali):

which means it's ?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

you cant do it.. so then it would be A

OpenStudy (nurali):

correct please click best response.

OpenStudy (anonymous):

thank you! :)

OpenStudy (nurali):

My Pleasure.